### v1

# Domain Adaptation-Enhanced Searchlight: Enabling brain decoding from visual perception to mental imagery

Abstract
In cognitive neuroscience and brain-computer interface research, accurately predicting imagined stimuli is crucial. This study investigates the effectiveness of Domain Adaptation (DA) in enhancing imagery prediction using primarily visual data from fMRI scans of 18 subjects. Initially, we train a baseline model on visual stimuli to predict imagined stimuli, utilizing data from 14 brain regions. We then develop several models to improve imagery prediction, comparing different DA methods. Our results demonstrate that DA significantly enhances imagery prediction, especially with the Regular Transfer approach. We then conduct a DA-enhanced searchlight analysis using Regular Transfer, followed by permutation-based statistical tests to identify brain regions where imagery decoding is consistently above chance across subjects. Our DA-enhanced searchlight predicts imagery contents in a highly distributed set of brain regions, including the visual cortex and the frontoparietal cortex, thereby outperforming standard cross-domain classification methods. The complete code and data for this paper have been made openly available for the use of the scientific community.
keywords:
Domain Adaptation , Brain Decoding , fMRI
journal: Neurocomputing
[…]
[inst1]organization=Intelligent Systems Group at the University of the Basque Country (UPV/EHU),city=San Sebastián, country=Spain
[…]
1 Introduction
Mapping brain activation patterns to their corresponding cognitive states is known as brain decoding. Methods for brain decoding are instrumental for understanding the brain processes involved in cognition [1] and therefore, an increasing number of works address the question of solving brain decoding tasks with machine learning algorithms [2, 3, 4].
Specifically, the degree to which visual imagery depends on the same neural mechanisms as visual perception remains a subject of ongoing research. Previous studies using standard multivoxel pattern classification analyses (MVPA) showed that a classifier trained to discriminate the contents of perception in the visual cortex can predict the contents of visual imagery [5, 6, 7, 8]. However, evidence that perception and imagery share neural representation in frontoparietal areas is scarce [9, 10].
In the present study we hypothesize that failures to reveal similar neural representations may be due to the classification pipeline not accounting for the different distributions of the data across perception and imagery domains. Consequently, we claim that Domain Adaptation (DA) [11, 12] methods are an adequate approach to alleviate the hypothesized distribution shift and enhance the transferability of classifiers trained in visual perception to the imagery testing data.
Thus, we investigate the impact of different DA approaches in a particular fMRI cross-domain experiment where subjects were asked to solve perceptual and imagery tasks. Cross-domain scenarios involving imagery are the basis for Brain-Computer Interfaces (BCIs) [13] and have been addressed before [6, 14, 7, 15] but, to the best of our knowledge, the literature is scarce when it comes to applying DA techniques to such problems using fMRI data. Furthermore, rather than restricting our analysis to the visual cortex, we explore the whole brain using one local neighbourhood classifier per voxel (a searchlight procedure, as introduced by [16]). In this procedure, the brain is partitioned into thousands of overlapping spheres, and the voxels from each sphere are assigned to a classifier. The searchlight method combats the curse of dimensionality by limiting the radius of the sphere, and offers an anatomically interpretable output.
In this paper, we integrate a DA technique into the searchlight method to unveil common neural representations for perception and imagery across the brain.
In summary, the contributions made in this paper are:
  1. 1.
    First extensive evaluation and comparison of well-established DA techniques on cross-domain classification problems defined on fMRI data, considering several brain Regions Of Interest (ROIs).
  2. 2.
    Introducing the DA-enhanced searchlight approach using the best-performing DA technique on the whole brain, to gain more understanding on the nature and location of the domain-shift.
2 Methodological background
We define a domain $D$ as a combination of input and output spaces and an associated probability distribution: $(X,Y,p_{D})$. In the simplest case, there is a source domain $S$ with enough data to build a reliable learner, and we aim to generalize to a different target domain $T$ for which there is limited knowledge. In this paper, we consider the situation in which both domains share input and output spaces, meaning that they have the same dimensionality and live in the same space, but data come from different probability distributions $p_{S}$ and $p_{T}$
From the ML point of view, the problem we are interested in is risk minimization in the target domain, given a set of data $(x_{i},y_{i})\sim p_{S}(x,y)$ from the source distribution (following notation from [17]). Although the formulation is valid for a general output space $Y$, we will restrict ourselves to binary classification. In this setting, we would like to find the best learner $h^{\star}$, defined as the one with minimum expected loss $l$ under the target distribution, $h^{\star}=arg\underset{h\in\mathcal{H}}{min}R_{T}(h)$, where $h$ is any learner included in a given hypothesis space $\mathcal{H}$. The expression for the risk includes a term $p_{T}(x_{i},y_{i})$, that represents the probability of the source samples under the target distribution (Eq. 1, left hand side).
Introducing into the integral the probability of the observations coming from their actual source distribution $p_{S}$, we arrive to an expectation under $p_{S}$ of the loss multiplied by the ratio of target to source distributions (Eq. 1 right hand side).
[…]
In practice, this expectation can be estimated using the sample average (empirical risk minimization) [18]. Depending on the structure of $p_{S}$ and $p_{T}$, the quantity $p_{T}(x,y)/p_{S}(x,y)$ can be further simplified by factoring into conditional and marginal distributions, giving rise to special cases of distribution shifts. For example, if the prior probabilities of $y$ differ between domains, but the conditionals $p(x|y)$ are close enough, we have a case of prior shift. Other particular settings include covariate shift (similar posterior distributions $p(y|x)$ for $T$ and $S$) and concept shift (similar data distributions $p(x)$ but $p_{T}(y|x)\neq p_{S}(y|x)$). In general, though, the data-set shifts have a more complicated and unknown structure. Depending on the nature of the distribution shift, certain DA approaches may be more suitable than others [19].
DA requires some previous knowledge of the target domain, be it in the form of some labeled (supervised DA, also known as few-shot learning) or unlabeled (unsupervised DA) instances from the target distribution. The amount of target domain knowledge incorporated into a DA technique is usually too scarce to build a model that relies in that information alone. In fact, for most DA applications, acquiring (or labeling) data from the target domain is difficult and/or expensive, but there is enough source domain data to build and validate a model with an adequate performance of unseen samples from that same source domain.
Given that such a model is achieved, generalization to novel examples from the target distribution can be tackled by several strategies, namely (a) instance-based, (b) feature-based and (c) parameter-based approaches, as summarized in Table 1. The naïve inclusion of the available target domain knowledge as additional training samples can also be considered as a rudimentary form of DA.
Main approaches to Domain Adaptation
Methods and key ideas Examples
Supervised Unsupervised
1. Instance-based 1. Measure the difference between each source domain and target domain instances. 2. Adjust the domain instance contributions during training. 1. Real unequal eigenvalues of same sign 1. Balanced Weighting [20] 2. TrAdaBoost [21] 3. IWN [22] 1. Real unequal eigenvalues of same sign 1. KMM [23] 2. ULSIF [24] 3. RULSIF [25]
2. Feature-based 1. Look for a space of common features with respect to the task on source and target domain. 2. Build a representation under which $p_{S}$ and $p_{T}$ are similar. 2. Feature-based 1. FA and PRED [26] 2. Real unequal eigenvalues of opposite sign 1. SA [27]
3. Parameter-based 1. Adapt the parameters of a pre-trained source-only model to build a suitable model with the same structure on the target domain. 3. Equal eigenvalues Subtype 1: Two Independent vectors 1. Regular Transfer [28]
Table 1: Summary of main approaches to Domain Adaptation
3 Related Work
In this section we review a number of works related to our research. We cover: 1) ML techniques for investigating commonalities in brain mechanisms for visual and imagery tasks; 2) DA methods applied to other neuroscience scenarios.
3.1 ML techniques for investigating commonalities in brain mechanisms for visual and imagery tasks
There have been numerous attempts to elucidate whether visual perception and imagery arise from similar brain mechanisms. Since the development of brain imaging, the most prominent ML methods used to study this problem are multivariate pattern analysis (MVPA) and deep learning approaches.
MVPA considers patterns of activity across multiple brain regions simultaneously, as opposed to univariate analysis which looks at each voxel independently. It is a widespread method for testing cognitive theories [29].
When it comes to the relatedness of perception and imagery, all published work in the decade of 2010 involves attempts of pure transfer using MVPA [6, 14, 30, 7]. These pioneering works evaluate a source-only classifier in the data from the target domain without implementing DA techniques. As we have already discussed, pure transfer is not guaranteed to work, due to a distribution shift between the brain patterns induced by different experimental conditions. Indeed, all of the aforementioned papers suffer from low generalization, showing a notable decrease between classifier performance in perception data and the attempted classification of imagery data.
With a slightly different approach, [8] trained a voxel-wise encoding on visual low level features and then applied it to imagery brain patterns recorded while the participants were remembering a previously memorized work of art. Their model was able to identify the target artwork among a large dataset of options, supporting the idea that low level features are similar between perception and imagery. In 2017, [31] used several feature extraction methods trained on visual fMRI data to construct a generic decoder for both vision and imagery. They attained around 90% accuracy for seen object and around 60% accuracy for imagined object identification, revealing a significant performance gap. A review [32] from 2019 concluded that visual imagery and perception share a variety of neural mechanisms in the visual, parietal, and frontal cortex, but that those similarities are modulated by imagery vividness, and that temporal dynamics are different (as studied in 2020, see [33]).
In 2020 [10], MVPA was used to decode stimulus identity in visual imagery of simple shapes and letters, identifying the regions where such decoding was possible. A searchlight analysis revealed cross-domain decoding in the occipital, parietal and inferotemporal cortices.
From 2023 onwards, deep learning methods have been employed to reconstruct mental images from brain activity patterns [34, 35], study visual object representation with a high spatiotemporal resolution [36] or learn robust representations from EEG recordings [37]. In [38], authors conduct a representational similarity analysis comparing the parameters of a CNN trained to recognize digits and a fMRI data-set containing brain patterns linked to digit recognition.
However, as pointed out by the authors in [32], the fact that it is possible to have imagery deficits (e.g. aphantasia) without associated perceptual issues indicates that the simulation processes underlying perception and imagery are not exactly the same. More evidence of these differences was found in [39], where hierarchical generative networks reveal distinct features for visual perception and imagery in low-level visual areas.
In this work, we claim that DA techniques are an additional tool to investigate and quantify the degree of difference between brain mechanisms associated with perception and imagery, as well as to mitigate the cross-domain brain decoding performance gap shown in [14, 6, 40]. Firstly, we consider several ROIs to determine the best DA approach for our use case. Then, by conducting a searchlight analysis across the whole brain, we investigate which brain regions differ the most between both domains, and improve the classifier’s performance on a local level. The study of a large number of subjects offers some insight into the interpersonal variability of the hypothesized domain-shift, as well as the possibility to discover brain regions where such hypothesized domain-shift occurs on a consistent basis.
3.2 Other applications of DA in neuroscience
DA has been used in other neuroscientific settings, mostly in cross-device experiments for medical diagnosis. For instance, in [41], authors propose a low-rank DA method for autism diagnosis, aiming to determine a common lower-dimensional representation for data from multiple imaging devices. A very different approach involves the field of optimal transport, in which DA has been used to align the distributions of fMRI images from multiple devices with the goal of brain dementia identification [42]. Both of these methods fit into the category of feature-based DA, because they entail the projection of source and target domains to a new feature space where the domain shift has been minimized.
The use of instance-based approaches has also been reported in the literature. For example, in [43], the authors propose a re-weighting scheme for the diagnosis of Alzheimer’s disease. In the field of brain tumor segmentation, instance-based DA has been employed to correct sample selection bias amongst MRI image patches [44]. In [45], the authors compare the performance of four DA techniques in a cross-device MRI image segmentation task using support vector machines. Specifically, they review three instance-based approaches (including TrAdaboost [21]) and a parameter-based technique, similar to Regular Transfer [28].
Regarding other parameter-based methods, most published works deal with fine-tuning of deep learning models. For example, authors of [46] study the benefits of fine-tuning in a brain lesion segmentation task, adapting CNN models pre-trained on brain MRI scans. Some other works leverage state-of-the-art computer vision models trained in large-scale multipurpose data-sets, adapting the final layers to their task [47, 48]. Beyond neuroscience, more examples of DA in medical imaging are referenced in [49].
A recent survey [50] advocates for including DA in the computational neuroscientist’s toolkit, highlighting the need to test state-of-the-art DA methods in scenarios of complex data such as neuroimaging.
In summary, many neuroscience questions have benefited from DA, but there is a literature gap when it comes to cross-domain classification. Additionally, most published works in the intersection of neuroscience and DA are focused on developing new approaches for very specific tasks and using only one or a reduced set of DA approaches, without extensive evaluation of standard DA methods. In contrast, we conduct a thorough comparison of existing DA methods for a cross-domain task in a large number of independent data-sets, with the goal of quantifying the domain-shift between visual perception and imagery.
4 Description of the cross-domain analysis
4.1 Experimental setup
We quantified the contribution of 11 different DA algorithms to a binary cognitive state classification task using 18 data-sets, each containing around 9000-14000 anatomically selected voxels from approximately 500-600 fMRI images of a specific subject gathered in independent realizations of the same experiment, described in detail in [40]. The voxels come from 14 ROIs, the exact dimensionality of which is subject-dependent (Suppplementary Table 3). The ROIs were extracted using FreeSurfer [51], and their names are listed in Supplementary Table 2.
The experiment from which the data-sets were collected consisted on two phases: (1) perception and (2) imagery. In the perception step, each participant was presented with living (Dog) and non-living (Scissor) images from the THINGS [52] database. In the imagery stage, they received auditory cues instructing them to vividly imagine an object from one of those two categories for several seconds. The goal was to classify imagery instances into living or non-living, leveraging a classifier trained mostly on perception instances.
In the perception phase, subjects participated in 4 runs of scanning. Each run was composed of 14 trials, 7 for each category (living/non-living). In each trial, participants observed a sequence of 12 images belonging to the same category. There was an 8.5-second rest period between trials, to ensure separation between brain activity patterns associated with each category. The imagery phase followed an analogous structure but consisted of only 2 runs.
4.2 Machine learning task and validation framework
Taking the delay of the hemodynamic response and the timing of stimulus presentation into account, there was not a one-to-one correspondence between stimuli and fMRI images within each trial, but those from different trials were safely considered independent. Thus, in all our analysis, we avoided data leakage ensuring that all instances of the same trial belonged either to the train or the test set.
We used 100 different partitions of the data to validate our methodology. Each partition determined the exact assignment of instances into train or test sets. The source domain (perception) train/test ratio was 4:1, with approximate class stratification and observing the trial constraint. As for the target domain, we randomly selected $N_{t}$ instances with approximate class stratification to compose the training set, discarded the rest of the instances belonging to those trials, and used the remainder for testing. $N_{t}$ is a parameter of the methodology, which we varied from 10 to 100 in steps of 10.
For each train/test split, we fitted a baseline estimator $h$ on the source training set, as well as another instance of the same standard classifier on the union of source and target training sets ($h_{NAIVE}$). We evaluated their performance on the target test set. $h$ is considered to be the experiment baseline, and $h_{NAIVE}$ accounts for the effect of naïvely including some limited knowledge of the target domain into a standard classifier without considering DA techniques. To assess the contribution of a specific DA technique, we fitted a new instance of the chosen estimator on the source domain training set $(X_{s},y_{s})$ alone and incorporated the target training instances $(X_{t},y_{t})$ through said DA technique, obtaining a new estimator, $h_{DA}$, which we evaluated on the test sets of the target domain. We chose Logistic Regression as an estimator due to its simplicity, computational speed and compatibility with all DA methods under study.
For each subject, we compared the performance of 11 DA approaches on the union of all ROIs using the scikit-learn compatible implementations in the ADAPT [53] library. B contains a brief description of each method.
As mentioned above, in the first place we evaluated the contribution of each DA technique to our task in the different subjects using the voxels from all 14 ROIs. Then, based on the comparison results, we also evaluated the performance of the best technique beyond the 14 ROIS, considering all the voxels from the whole brain. This whole-brain analysis was conducted using the searchlight [16] procedure.
4.3 Evaluation metrics and statistical validation
4.3.1 Initial comparison of DA techniques
The evaluation criteria on which our comparisons are based are detailed hereafter. First, note that the methodology was constructed in such a way that the training of $h$, $h_{NAIVE}$ and $\{h_{DA}\}_{DA=1}^{11}$ were performed on exactly the same data, and the same is true for the evaluation phase of all estimators. During the 100 repetitions of the process, the balanced accuracies of each of the 13 estimators were measured on the test set of the target domain, obtaining the 13 corresponding performance vectors of size 100. Those 13 balanced accuracy vectors of $h$, $h_{NAIVE}$ and $\{h_{DA}\}_{DA=1}^{11}$ can be thought of as paired observations.
To quantify the amount of information needed to actually profit from DA methods, we applied our methodology with a varying amount of target domain instances $N_{t}$ between 10 and 100 in steps of 10, conducting the same comparison. By doing that, we determined the effect of incorporating a growing amount of target training instances into both $h_{NAIVE}$ and $h_{DA}$; that is, whether $h_{DA}$ profits from this information more than $h_{NAIVE}$ does.
Hence, our results are in the form of 10 balanced accuracy tables per subject, one corresponding to each value of $N_{t}$. Each table has 13 columns (one per ML approach) and 100 rows (one per data partition).
Firstly, we concatenated the 10 tables pertaining to each subject and conducted an all versus all statistical comparison based on the Friedman aligned ranks test, with a Shaffer correction for multiple comparisons. The non-parametric character of the Friedman test avoids distributional assumptions on our results, and the Shaffer correction controls the family-wise error rate with a low computational cost compared to other correction methods.
We also concatenated all of the results (i.e. each of the 180 tables pertaining to each subject and value of $N_{t}$) to obtain a global comparison between all methods.
All the statistical analyses detailed in this section were based on the implementations in the R library scmamp [54].
4.3.2 Evaluation of the searchlight analysis
For the searchlight analysis, we used the same data partitioning methodology as for the ROI-based analysis. In this case, each sample contains a whole-brain three-dimensional image with around 40,000 voxels (the exact number of which is subject-dependent - see Supplementary Table 3). Each voxel is a cube with a 3mm side, and we defined the radius of the searchlight spheres to be 12mm.
We fitted three independent Logistic Regression classifiers on each sphere:
  1. 1.
    $h$: Trained only on the samples from the source domain to establish a baseline.
  2. 2.
    $h_{NAIVE}$: Trained on the union of samples from the source domain and 100 samples from the target domain, to assess the effect of including imagery data without DA.
  3. 3.
    $h_{RTLC}$: Incorporating the same 100 samples from the target domain into $h$ via the RTLC algorithm.
All classifiers were tested on the same holdout instances from the target domain. $h$ was also tested on a hold-out set from the source domain. The procedure was repeated for 100 different data partitions, and the mean balanced accuracy per sphere was computed, obtaining a three-dimensional brain map of performance scores for each subject and classifier. To carry out these analyses, we integrated some functionality from ADAPT into the searchlight tools from the Nilearn library [55].
Then, we substracted 0.5 (the chance level balanced accuracy) from all the maps for further statistical testing.
To be able to extract meaningful anatomical insights and make comparisons between subjects, each score map was transformed to the standard Montreal Neurological Institute (MNI) space using the software FSL [56]. Subsequently, the maps from all the subjects were merged into a single four-dimensional image.
The statistical validation of our results employed the randomise technique [57] as implemented in FSL. This permutation-based strategy requires only minimal assumptions for validity (in our use case, exchangeability of the subjects and a symmetric distribution of the centered maps at a voxel level) and accounts for multiple comparisons when testing voxel by voxel [58]. Using this approach, we conducted a one sample t-test with variance smoothing ($\sigma=6$mm) and 10000 permutations. Threshold-Free Cluster Enhancement [59] was used, and the the family-wise error rate was controlled. This procedure yields a three-dimensional brain map of p-values, where each voxel indicates the probability that the balanced accuracy measured at said voxel across different subjects is statistically greater than chance. We chose a significance threshold of $\alpha=0.05$.
We also used randomise to detect above-chance decoding for each subject independently, identifying informative clusters on an individualized basis. Although [58] advises against using this method on time-correlated data from the same subject, we can safely use it here because we are permuting accuracy values obtained in different data partitions with no time-series structure.
5 Code and data availability
The code for this paper is available at the github repository AlexOlza/DA-enhanced-searchlight. The data is available at the Open Science Framework (OSF) platform at https://osf.io/audmx/.
6 Experiments
6.1 Comparison of DA techniques
Figure 1 shows the unique Critical Difference (CD) diagram obtained by concatenating all the observations (both across $N_{t}$ and across different subjects). The scale above shows the average ranking of each algorithm across all observations, and the horizontal lines group together algorithms that are not statistically different. From this, we conclude that RTLC was the best algorithm, followed by the group of NAIVE, TAB, BW and FA which are statistically indistinguishable ($\alpha=0.05$). On the other hand, SA is shown to be the worst.
Figure 1: Critical Difference diagram. The scale above shows the average ranking of each algorithm across all observations, and the horizontal lines group together algorithms that are not statistically different.
Figure 2 summarizes the results of the tests in the 18 CD diagrams arising from concatenating the 10 tables corresponding to different $N_{t}$ values for each subject, which are shown in Supplementary Figure 9. For each cell $(i,j)$ in this table, we report the number of subjects for which algorithm $j$ was significantly better than algorithm $i$ ($\alpha=0.05$). This can be extracted from the CD diagrams by counting how many times $j$ appears to the left of $i$ and substracting the number of times they are joined by a horizontal line. The sum of each column represents the number of times each algorithm was better than others and, complementarily, the sum of each row reports the number of times each algorithm was significantly worse than others.
Figure 2: Frequency table. For each cell $(i,j)$, the numerical value shows for how many subjects algorithm $j$ was significantly superior to algorithm $i$. The sum of each column accounts for the number of times each algorithm was significantly better than others, and the sum of each row shows how many times each algorithm was significantly worse than others.
From Figure 2, we extract the following highlights:
  • 1.
    RTLC outperformed others a total of 127 times, and was outperformed 51 times. It was the most outperforming and least outperformed algorithm.
  • 2.
    RTLC was better than BASELINE for 11 subjects, worse than BASELINE for 4 and statistically the same for the remaining 3.
  • 3.
    RTLC was better than NAIVE for 10 subjects, worse than NAIVE for 4 and statistically the same for the remaining 4.
  • 4.
    NAIVE was better than BASELINE for 10 subjects and worse for 8.
  • 5.
    SA was the most outperformed and least outperforming algorithm.
6.2 Analysis of individual ROIs using RTLC
Figure 3 shows the CD diagram of the RTLC performance in the imagery domain across individual ROIs, revealing the most informative regions for this method to be the fusiform gyrus (FFG), the inferior parietal lobe (IPL), the inferior temporal gyrus (ITG) and the middle temporal gyrus (MTG).
Similarly, Figure 4 highlights the fusiform gyrus as the most informative region in perception according to the baseline, with an average ranking of 1, meaning that it was the most informative region for all subjects and data partitions.
For completeness, Figure 5 shows the ranking of ROIs acording to the baseline in the imagery domain. Although the fusiform gyrus remains the most informative ROI in this domain, its average ranking has increased, indicating subject variability in the imagery decoding across individual ROIs.
Figure 3: CD diagram of RTLC performance in the imagery domain across individual ROIs. FFG: Fusiform Gyrus; LOG: Lateral occipital gyrus; ITG: Inferior temporal gyrus; TP: Temporal pole; PCG: Posterior cyngulate gyrus; PCUN: Precuneus; IPL: Inferior parietal lobe; MTG: Middle temporal gyrus; SFG: Superior frontal gyrus; IFGoperc: Inferior frontal gyrus, pars opercularis, IFGtriang: Inferior frontal gyrus, pars triangularis; IFGorbital: Inferior frontal gyrus, pars orbitalis; FP: Frontopolar cortex; MOG: Medial orbital gyrus.
Figure 4: CD diagram of baseline performance in the perception domain across individual ROIs. FFG: Fusiform Gyrus; LOG: Lateral occipital gyrus; ITG: Inferior temporal gyrus; TP: Temporal pole; PCG: Posterior cyngulate gyrus; PCUN: Precuneus; IPL: Inferior parietal lobe; MTG: Middle temporal gyrus; SFG: Superior frontal gyrus; IFGoperc: Inferior frontal gyrus, pars opercularis, IFGtriang: Inferior frontal gyrus, pars triangularis; IFGorbital: Inferior frontal gyrus, pars orbitalis; FP: Frontopolar cortex; MOG: Medial orbital gyrus.
Figure 5: CD diagram of baseline performance in the imagery domain across individual ROIs. FFG: Fusiform Gyrus; LOG: Lateral occipital gyrus; ITG: Inferior temporal gyrus; TP: Temporal pole; PCG: Posterior cyngulate gyrus; PCUN: Precuneus; IPL: Inferior parietal lobe; MTG: Middle temporal gyrus; SFG: Superior frontal gyrus; IFGoperc: Inferior frontal gyrus, pars opercularis, IFGtriang: Inferior frontal gyrus, pars triangularis; IFGorbital: Inferior frontal gyrus, pars orbitalis; FP: Frontopolar cortex; MOG: Medial orbital gyrus.
6.3 DA-enhanced searchlight analysis
Figure 6 shows the projection of the average balanced accuracy of each approach (computed across the subjects) onto the areas of statistically significant brain decoding. We see that only within-domain prediction attains good balanced accuracy when averaged across all the subjects, and that in this case (perception) decoding is possible in a highly distributed set of brain regions. For the baseline and naïve approaches, decoding is only possible in a small isolated region. On the contrary, RLTC is able to obtain significantly above chance imagery decoding in an extensively spread brain volume.
Figure 7 shows the results of the statistical comparison between RTLC and the naïve approach, highlighting only the statistically significant clusters ($p<0.05$, whole-brain corrected). The naïve method was never statistically significantly above the baseline.
Figure 8 accounts for the inter-subject variability of RTLC performance. Each set of brain slices corresponds to a different subject, and shows the average accuracy across the 100 data partitions projected onto the space where such average is statistically greater than chance level. Supplementary Figures 10, 11 and 12, respectively, show the subject variability of the perception decoding, the baseline cross-domain decoding and the cross-domain decoding with naïve inclusion of target instances in the training set.
Figure 6: Between-subject average of the balanced accuracy of the different approaches in the regions that obtained statistically above chance decoding.
Figure 7: Brain regions where RTLC decoding is significantly better than the naïve approach ($p<0.05$, whole-brain corrected).
Figure 8: Inter-subject variability of the DA-enhaced searchlight accuracy. For each subject, the color of each voxel represents the average RTLC accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically above 0.5.
7 Discussion
Although many neuroscience studies face challenges due to distribution shifts, most current analyses of cross-domain experiments in cognitive neuroscience do not address this issue. This oversight hinders the discovery of shared neural representations and limits the applicability of classifiers trained on source domain data to target domains. For instance, in decoded neurofeedback studies, the aim is to induce a specific activation pattern in the brain, usually without participant awareness of the goal state [60]. This procedure involves decoding the participant’s cognitive state in real time using a previously trained classifier. Then, a measure of the similarity between the current brain state and the goal activation pattern is computed, and the degree of similarity is presented as feedback to the subject. Several decoded neurofeedback works involve imagery decoding based on a stimulus presentation [40, 61] or motor task [62, 63], but none of them implement DA techniques to boost feedback accuracy.
We hypothesized that predictive models trained with fMRI data from different experimental conditions could benefit from domain adaptation (DA). Our results support this hypothesis, showing that DA using the RTLC technique significantly improves cross-domain brain decoding from perception to imagery compared to the naive inclusion of imagery instances in the training set. Furthermore, 6 out of the 11 DA methods studied outperformed the baseline (see Figure 1). With our DA-enhanced searchlight approach, we demonstrated statistically significant pattern similarities between perception and imagery in the visual cortex and beyond, including a distributed set of brain regions in the frontoparietal cortex.
Assessing the effect of DA on predictive ability offers a more complete perspective on the role of imagery versus perception in the brain representation of specific concepts. If DA is able to improve classification, it indicates that the initial differences in the distributions corresponding to each stimulus are sufficiently marked to prevent the baseline model from exploiting all the information contained in the data. Furthermore, when the model has a goal beyond testing hypotheses about neural activity, predictive improvement may be critical to the applicability of the classifier, for instance in neurofeedback protocols for clinical interventions.
Our results show that RTLC can be beneficial for cross-domain brain decoding when employed in the union of the ROIs under study, and also when integrated into a searchlight procedure. These results align with previous work stating that BCIs can be improved by incorporating DA into their training methodology [64, 65]. This requires little extra effort, aside from gathering a small sample of data from the target domain. In the case of fMRI studies, acquiring this data involves no significant additional cost. Importantly, we have shown that naïvely concatenating this data into the training set of a conventional classifier is less effective than using RTLC to incorporate the same information.
In the analysis across individual ROIs presented in Section 6.2, we assess the contribution of the different ROIs to RTLC performance, as well as their informativeness according to the baseline. We see that RTLC performs best in the FFG, but is also able to extract information from ROIs that are indirectly related to vision such as the inferior parietal lobe (IPL), the inferior temporal gyrus (ITG), and the middle temporal gyrus (MTG).
The success of DA is strongly limited by the signal-to-noise ratio of the data. When the task of interest is difficult in the source domain itself, resulting in unreliable starting models, generalization to the target domain will be poor regardless of the DA technique employed. In our case, not all ROIs contain valuable information on visual object category, which becomes evident in Figure 4 and in the searchlight analysis. Specifically, Figure 4 highlights the fusiform gyrus (FFG) as the most informative ROI for visual perception decoding in all cases, whereas ROIs such as the inferior frontal gyrus, pars orbitalis (IFGorbital), the frontopolar cortex (FP) and the superior frontal gyrus (SFG), among others, encode little information about visual object category. Equivalently, from Supplementary Figure 10 we conclude that all subjects have good perception-to-perception decoding in the areas of the occipital lobe associated with vision, but only some subjects encode visual information in additional brain regions. For around half of the subjects, the visual information is distributed on the whole brain above chance level, even though the decoding is always stronger in the occipital lobe. Hence, the between subject average decoding is predominant in the occipital lobe but also extends to other areas (Figure 6 - perception).
When it comes to the direct applicability of standard perception-based models to the imagery domain, Figure 6 (baseline and naïve) reveals a poor transferability on average. Inspecting Supplementary Figures 11 and 12, we discover that the extent of the domain shift (indicated by the performance gap of these approaches compared to perception prediction) is in fact very subject-dependent. Participants such as Subject 7 and Subject 14 exhibit good transferability, but others such as Subject 5 or Subject 8 have very sparse above-chance decoding regions in imagery. Crucially, we reiterate that the across-subject average of the decoding is limited to a very small brain region, which makes the standard models unreliable for additional out-of-sample subjects in the imagery domain.
In contrast, even when averaged across the subjects, DA-enhanced models attain above chance decoding in mostly the same areas as perception prediction does, although, expectedly, the acuraccy drops considerably. Once again, this is explained by the subject variability depicted in Figure 8. Some subjects benefit from DA more than others do. Participants such as Subject 8 remain a challenge, whereas for others like Subject 5 the adapted models are well tailored to the imagery scenario highlighting imagery-specific brain regions. In cases such as Subject 5, Subject 13 and Subject 10 the parietal lobe appears to play no role on encoding visual information that distinguishes between the living and non-living classes (see Figure 10), but it contains relevant imagery information which is discriminative of those classes.
Regarding why RTLC outperforms other DA methods in the ROI-based comparison, the simplicity of the technique might make it adequate to deal with high-dimensional data. In our comparison of DA techniques, the amount of examples is low compared to the amount of voxels (see Table 3 for the exact numbers). This may be detrimental to more sophisticated DA techniques, but still feasible for RTLC. In addition, RTLC is tailored to the specific estimator $h$ used for the source domain. Thus, if $h$ has good performance, it will be easier to obtain a good $h_{RTLC}$. This is evidenced in Supplementary Figure 10 and in Figure 8, which respectively indicate the performance of the searchlight $h$ and $h_{RTLC}$.
In fact, from the point of view of ML, it is important to analyze the limitations of current DA techniques in real-word scenarios. Usually, DA methods are tested considering artificial benchmarks such as Office-31 [66] and MNIST [67] or real problems with a low dimensionality. fMRI data offers the opportunity to evaluate these methods, elucidating whether the complexity of the methods is translated into an enhanced performance. Therefore, one of the goals of this paper is to evaluate the performance of state of the art DA techniques when dealing with high-dimensional, noisy and complex data.
From a more practical perspective, establishing an accurate learner that is capable to generalize from perception to imagery is relevant for many neuroscience applications, including but not limited to BCIs, neurofeedback training, and clinical scenarios. For instance, in rehabilitation settings, decoding brain activity patterns associated with mental imagery and visual perception could be used to develop interventions for individuals with visual impairments or neurological disorders affecting visual processing.
Among the limitations of our work, we lack a comprehensive characterization of the participants, which was not possible to achieve with the available data. Our results reveal a large amount of subject variability both in terms of the performance of baseline classifiers and the success of DA efforts, which is natural given the individuality of brain processes, but which we are not able to adequately explain. For instance, measuring the subjective vividness of visual imagery as done in [68] would have allowed us to establish the relationship between individual imagery strategies and the degree of improvement brought by DA. Another limitation is that some DA methods that did not excel at the ROI-level comparison could have been even better than RTLC in the searchlight analysis, which does not suffer from the curse of dimensionality. However, due to the computationally intensive nature of the searchlight procedure, a comparison of all the DA techniques on a voxel level would have been unattainable. This work provides a novel precedent of fMRI DA-enhanced searchlight for cross-domain brain decoding, while acknowledging that RTLC is not necessarily the best performing DA technique in this context. Nevertheless, sophisticated DA techniques and/or underlying estimators such as neural networks are impractical to integrate into a searchlight procedure that requires fitting tens of thousands of such models for each data partition.
The idea of a DA-enhanced searchlight could also be applied to other domain-shift scenarios. Future research lines include testing the validity of this approach in cross-subject and cross-device problems, where the input space is not shared across domains. This additional difficulty could be addressed with proper representation learning techniques.
8 Conclusions
In this paper, we have introduced an effective way to integrate DA into the searchlight procedure, significantly enhancing classification performance in cross-domain brain decoding. Through a comprehensive analysis involving data from 14 ROIs, we demonstrated that several DA techniques surpassed our baseline model, which was trained exclusively on perception data and evaluated on imagery data. This underscores the presence of a domain shift between the neural activities associated with visual perception and imagery of living versus non-living objects.
The best performing method was RTLC. Utilizing RTLC, we assessed both the nature and location of the domain shift within our DA-enhanced searchlight approach. Remarkably, RTLC enabled above-chance decoding of imagined stimuli in a highly distributed set of brain regions, including the visual cortex and the frontoparietal cortex. This contrasts sharply with non-DA searchlight procedures, which did not achieve comparable performance.
These findings are particularly promising for applications in BCIs and neurofeedback training, suggesting that DA techniques like RTLC can substantially improve the decoding of imagined stimuli or any other critical brain state that investigators may seek to induce in decoded neurofeedback studies. This advancement paves the way for more effective and reliable cross-domain brain decoding methods, potentially revolutionizing how neural data is utilized in practical applications.
9 Declaration of Competing Interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
10 CRediT author contributions
Olza: Conceptualization, Investigation, Software, Writing - Original draft Soto: Conceptualization, Data curation, Writing - review and editing, Funding acquisition, Supervision Santana: Conceptualization, Writing - review and editing, Funding acquisition, Supervision
11 Acknowledgements
This work has been supported by the Basque Government via the IKUR strategy, project IT1504-22 and the BERC 2022-2025 program, by the Spanish Ministry of Science and Innovation (PID2022-137442NB-I00), by Elkartek (KK-2023/00090), by the Spanish Ministry of Economy and Competitiveness, through the ’Severo Ochoa’ Programme for Centres/Units of Excellence (CEX2020-001010-S) and also from project grant PID2019-105494GB-I00.
Appendix A Supplementary Results
Abbreviation Name
FFG Fusiform gyrus
LOG Lateral occipital gyrus
ITG Inferior temporal gyrus
TP Temporal pole
PCG Posterior cingulate gyrus
PCUN Precuneus
IPL Inferior parietal lobe
MTG Middle temporal gyrus
SFG Superior frontal gyrus
IFGoperc Inferior frontal gyrus, pars opercularis
IFGorbital Inferior frontal gyrus, pars orbitalis
IFGtriang Inferior frontal gyrus, pars triangularis
FP Frontopolar cortex
MOG Medial orbital gyrus
Table 2: Abbreviations and names of the ROIs under study
Perception Imagery
Subject Voxels (ROIs) Voxels Instances Prevalence Instances Prevalence
1 10373 35744 368 0.503 234 0.513
2 10675 35464 362 0.500 233 0.494
3 10383 37775 362 0.494 232 0.509
4 10852 41513 362 0.517 233 0.511
5 10768 40194 362 0.506 233 0.494
6 10577 41162 364 0.495 233 0.498
7 11078 47814 364 0.500 231 0.511
8 11178 43787 366 0.511 233 0.502
9 13916 51748 361 0.504 234 0.491
10 11841 46884 362 0.497 231 0.506
11 10807 41465 362 0.503 233 0.511
12 12171 44995 364 0.500 232 0.496
13 12208 46920 365 0.496 233 0.506
14 10823 38441 365 0.496 234 0.500
15 12127 40111 360 0.500 234 0.491
16 9161 33991 364 0.489 232 0.500
17 12179 45649 363 0.510 233 0.511
18 11214 43522 361 0.507 234 0.504
Table 3: Dataset characteristics for each subject
Figure 9: Critical Difference diagrams for each subject, comparing different DA techniques on the union of 14 ROIs.
Figure 10: Subject variability of the within-domain searchlight accuracy. For each subject, the color of each voxel represents the average perception-to-perception accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically significantly above 0.5.
Figure 11: Subject variability of the cross-domain searchlight accuracy. For each subject, the color of each voxel represents the average baseline perception-to-imagery accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically significantly above 0.5.
Figure 12: Subject variability of the cross-domain searchlight accuracy with 100 target domain instances available during training. For each subject, the color of each voxel represents the average accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically significantly above 0.5.
Appendix B Brief description of each DA method used
  • 1.
    BalancedWeighting [20]: A supervised, instance-based method that fits the base estimator $h$ source and target labeled data according to the modified loss $(1-\gamma)\mathcal{L}(h(X_{s}),y_{s})+\gamma\mathcal{L}(h(X_{t}),y_{t})$, where the parameter $\gamma$ controls the ratio between fitting the source data and adapting to the target data, with higher values indicating more adaptation. In our case, $\mathcal{L}$ is the binary cross-entropy and $\gamma=0.5$. BalancedWeighting can be used with any supervised estimator.
  • 2.
[…]
  • 3.
    KMM (Kernel Mean Matching) [23]: An unsupervised, instance-based sample bias correction approach that minimizes the Maximum Mean Discrepancy (MMD) between source and target domains by solving a quadratic optimization problem such that the means of $X_{s}$ and $X_{t}$ in a reproducing kernel Hilbert space (RKHS) are closer than a predefined threshold. We used the default Radial Basis Function kernel.
  • 4.
    TrAdaBoost (Transfer AdaBoost for Classification) [21]: A supervised, instance-based iterative reweighing approach. Firstly, an initial estimator is fitted on the concatenation of labelled source and target data. Then, the samples are reweighed based on the performance of said estimator, decreasing the weight of the wrongly classified source instances under the hypothesis that those instances are more dissimilar to the ones coming from the target domain.
  • 5.
    FA [26]: A supervised, feature-based pre-processing step leveraging feature augmentation, which aims to separate features into (a) those specific to the source domain, (b) those specific to the target domain and (c) those common to both, without altering the model fitting stage, so it can be applied to any supervised estimator.
  • 6.
    PRED [26]: Another feature-augmentation DA technique, mixing the idea in FA with the notion of model stacking. First, the base estimator is trained solely on source data, predicting target values on the available target data. These predictions are concatenated with $X_{t}$ as a new feature, and a new instance of the estimator is trained on the resulting data-set to obtain the final $h_{DA}$.
  • 7.
    ULSIF (Unconstrained Least-Squares Importance Fitting) [24]: An unsupervised, instance-based method to correct the difference between input distributions of source and target domains as measured by their relative Pearson divergences by reweighing the source instances. This entails solving a quadratic optimization problem, which has a closed-form solution involving the inverse of a matrix $H$ of size equal to the number of samples in $X_{s}$. This matrix contains the element-wise products of a kernel transformation applied to $X_{s}$. $H^{-1}$ is then multiplied with the kernel transformation of $X_{t}$ to give the weights of the source samples. We used an RBF kernel.
  • 8.
    NearestNeighborsWeighting [69]: An unsupervised, instance-based approach that computes, for each source domain sample, the number of neighbours inside a certain radius in the target data set. Source instances are reweighted proportionately.
  • 9.
    RULSIF (Relative Unconstrained Least-Squares Importance Fitting) [25]: A very similar approach to ULSIF, but allowing the target instances to influence the matrix $H$ via an additive contribution of the element-wise products of their kernel transformation.
  • 10.
    SA (Subspace Alignment) [27]: An unsupervised, feature-based method that linearly aligns the source domain to the target domain in a reduced PCA subspace of a certain pre-specified dimension.
  • 11.
    IWN (Importance Weighting Network) [22]: An unsupervised, instance-based method that reweighs the source domain samples minimizing the MMD between the reweighed source and the target distributions by using a neural network to predict the sample weights, thus reducing the computational burden of KMM.